find the missing length of the right triangle
bottom is 17 and the bottom is 15

Answer by jerrybassmasta(14)

I am assuming that when you refer both the bottoms, you mean the legs.
This means that all we have to do is find the hypotenuse.
The formula for the Pythagorean Theorem is represented by a^2+b^2=c^2
A and B meaning the legs(bottoms) and C being the hypotenuse.
17^2+15^2=c^2
289+225=c^2
514=c^2
After squaring both numbers you get
c=22.6